### v1

# Challenges in Testing Large Language Model Based Software: A Faceted Taxonomy

Abstract.
Large Language Models (LLMs) and Multi-Agent LLMs (MALLMs) introduce non-determinism unlike traditional or machine learning software, requiring new approaches to verifying correctness beyond simple output comparisons or statistical accuracy over test datasets.
This paper presents a taxonomy for LLM test case design, informed by both the research literature, our experience, and open-source tools that represent the state of practice. We identify key variation points that impact test correctness and highlight open challenges that the research, industry, and open-source communities must address as LLMs become integral to software systems.
Our taxonomy defines four facets of LLM test case design, addressing ambiguity in both inputs and outputs while establishing best practices. It distinguishes variability in goals, the system under test, and inputs, and introduces two key oracle types: atomic and aggregated. Our mapping indicates that current tools insufficiently account for these variability points, highlighting the need for closer collaboration between academia and practitioners to improve the reliability and reproducibility of LLM testing.
Large Language Models, Software Testing, Correctness, Ambiguity, Aggregated Oracles
1. Introduction
Large Language Models (LLMs) and Multi-Agent LLMs (MALLMs) (Wang et al., 2024; He et al., 2024) are transforming software development—not just through their capabilities but also due to their inherent non-determinism. Unlike traditional systems, where variability arises from e.g. unreliable servers, stochastic sub-components, or real-time inaccuracies, LLMs exhibit fundamental unpredictability due to their construction, differences in model selection, configuration, and input variations at both syntactic and semantic levels. Conventional testing methods and approaches to oracle formulation and correctness (Barr et al., 2014) struggle in this setting. Based on our experience in designing, implementing, testing, and assessing several software systems/solutions either partly or fully built around LLMs (Feldt et al., 2023; Yoon et al., 2024; Kaarre et al., 2024; Khoee et al., 2024; Wang et al., 2025; Kang et al., 2023, 2024b, 2024a, 2025), we propose a taxonomy for LLM test case design that captures key challenges and nuances in this emerging paradigm.
Recent research highlights the impact of input and output ambiguity in LLM-based applications. Subtle prompt variations can invert model responses, even under high-confidence settings (Xiao et al., 2024), and repeated queries–despite deterministic configurations (e.g., temperature = 0)–can produce inconsistent outputs (Atil et al., 2024). This variability raises concerns for replicability and necessitates advances in automated oracle design (Molina et al., 2024).
Traditional software testing relies on deterministic oracles, but the probabilistic nature of LLMs challenges this assumption. While Barr et al. (Barr et al., 2014) introduced probabilistic oracles to handle non-determinism, their framework does not account for prompt-driven variability in LLMs and MALLMs. More broadly, ML testing requires a paradigm shift (Braiek and Khomh, 2020), yet existing work does not explicitly treat variation as a first-class concern. The challenge is further compounded by LLMs’ hybrid nature, where behavior emerges from a combination of code, model inference, and prompt engineering.
This divergence from traditional testing is evident in how correctness is defined. Existing paradigms–whether deterministic (Barr et al., 2014), stochastic (Grunske, 2008; Grunske and Zhang, 2009; Barr et al., 2014), or ML-specific (Riccio et al., 2020; Dobslaw and Feldt, 2023)–struggle to address the multi-layered complexity of LLM-based systems. Unlike static ML models, where uncertainty stems from training data and inference, LLMs introduce an additional axis: the prompt itself, which acts as both input specification and behavioral modifier. Moreover, LLM performance may degrade post-deployment due to data shifts, yet monitoring mechanisms remain underdeveloped (Braiek and Khomh, 2020). The disconnect between LLM testing tools and core ML or SE testing literature further underscores the need for specialized methodologies (Hudson et al., 2024).
To address this gap, Section 2 introduces our faceted taxonomy (Usman et al., 2017) for LLM test cases, categorizing key variation points that impact evaluation correctness, including the distinction between Atomic and Aggregated Oracles. Unlike prior work–(Xiao et al., 2024) on adversarial robustness and (Hudson et al., 2024) on broad LLM testing taxonomies–our framework organizes test case design across multiple facets, extending beyond specific testing foci and high-level categorization. Section 3 then maps current practices onto this taxonomy, assessing coverage across dimensions. Finally, Section 4 identifies key challenges for ensuring correctness in LLM-based software.
2. Taxonomy for LLM Test Case Design
Testing LLMs presents unique challenges due to their non-deterministic behavior and the ambiguity inherent in both inputs (e.g., prompt phrasing and intent) and outputs (e.g., diverse, context-dependent responses). Beyond variability, practical concerns such as cost and granularity in test case design further complicate reliable evaluation. To address these challenges, we propose a structured taxonomy for LLM testing, organized around four core dimensions: Software Under Test (SUT), Goal, Oracles, and Inputs. Each of these core dimensions is further sub-divided into finer-grained sub-facets, capturing the specific factors that influence test case design and evaluation as well as points of variability between specific test case runs. This taxonomy provides a systematic framework for developing and refining test cases across the software lifecycle, emphasizing the need for adaptive and continuous testing strategies tailored to the dynamic nature of LLM-based systems. The order of these dimensions reflects their logical dependencies in test case design.
  1. (1)
    SUT is presented first because it defines the underlying system or component to be evaluated. The SUT typically remains constant across multiple test cases, serving as the foundation upon which different testing scenarios are applied.
  2. (2)
    The test case Goal comes next, as it specifies the unique objective of the test case and highlights the specific properties of it to be tested. While the SUT may remain unchanged, each test case targets a specific property, such as ensuring functional correctness or non-functional goals like safety, fairness, or robustness.
  3. (3)
    Oracles are then defined for each property derived from the test case goal.
  4. (4)
    Inputs are the final dimension, consisting of the datasets, user interactions, or synthetic prompts used to elicit responses from the SUT, which are then assessed using the defined oracles.
This ordered structure can help ensure clarity in the design process. Starting with the SUT allows for reusable test infrastructures, while distinct goals enable focused evaluations. The oracles then provide concrete evaluation mechanisms for each property, and the inputs serve as the vehicle to generate diverse test executions. However, the taxonomy is not prescribing a particular order of analysis and developing and testing with LLMs is an inherently iterative process. Below we detail each dimension further and discuss its sub-facets.
2.1. Software Under Test
The system under test (SUT) refers to the software implementation being evaluated for correctness. In this context, we define the SUT as a software system that integrates one or more large language models (LLMs)–such as in multi-agent LLM architectures (Yoon et al., 2024) (e.g., MALLM)–to achieve the expected behavior specified in its design. To analyze variation both across test cases and within a single test execution, we break down the SUT in LLM-centric systems into the following key sub-facets.
The Component represents the specific functional unit or role within the system that is under test. In LLM-based systems, this often corresponds to a prompt (or prompt template), an instruction set, or an agent role. For instance, in a multi-agent setup, the component might be the Planner agent (as in the multi-agent LLM system of (Yoon et al., 2024)) or a specific prompt guiding a data transformation task. However, our taxonomy remains agnostic to the level of testing: in a multi-agent LLM system, a targeted subsystem composed of multiple agents may itself serve as the SUT. The component facet thus defines the logical function under test, independent of the specific model(s) executing it.
The Model(s) refers to the LLM(s) implementing the component(s) during test execution. Different base models (e.g., Claude 3.5, Deepseek R1, GPT-4o) can produce distinct behaviors even under identical prompts, making this a critical variation point. In MALLMs, each component may rely on a different model or a combination of models, requiring test cases to account for these permutations.
Each model operates under specific Configurations that directly influence its behavior. These include model parameters such as temperature, top-k/top-p sampling, and maximum token limits, as well as system-level settings like API rate limits and external tool integrations. Configurations may vary per model or be shared across components, introducing additional variability that must be managed in test design.
In the context of a specific test case, the SUT is a concrete instance with fixed choices for the Component, Model(s), and Configuration(s). However, our taxonomy explicitly distinguishes between fixed and variable elements across test runs. Any modification–such as swapping the underlying LLM, adjusting a temperature setting, or altering the role’s prompt–creates a new SUT instance that may behave differently under the same input.
This approach contrasts with traditional software testing, where the SUT is typically treated as fixed and configuration changes are minor or peripheral. In LLM-based systems, even small variations can significantly impact behavior, making them a deliberate factor in test case design rather than an unintended side effect. The complexity increases further in hybrid systems combining LLMs with conventional code, where changes in either can affect overall system behavior.
Traditional software testing relies on a specification that defines the expected behavior, interfaces, and constraints of the system under test (SUT). In LLM-based systems, this process becomes more complex due to the flexibility and ambiguity of natural language interactions. To manage this, developers often use prompt templates—structured inputs with placeholders, such as:
Write code to read <FORMAT> file in <PROGRAMMING_LANGUAGE> with telling variable names and add code comments.
These templates function like macros in traditional software, offering structured yet customizable input formats controlled by developers.
While prompt templates enable diverse inputs, they also complicate testing. LLMs typically generate free-text responses without strict adherence to requested formats, making automated evaluation challenging. To address this, invariant checks assess both syntactic correctness (e.g., ensuring the output is valid Python code) and semantic quality (e.g., verifying meaningful documentation). However, LLM responses can be ambiguous or only partially fulfil a request, requiring correctness evaluations to account for deviations ranging from minor formatting errors to complex, context-dependent variations.
These challenges, along with strategies for handling them, are further explored in the Oracle section. The Inputs section discusses how test cases instantiate prompt templates with varying, concrete data for evaluation.
2.2. Goal
The goal defines the high-level objective of a test case, which is then refined into specific, measurable sub-goals called Properties–the concrete conditions the SUT must satisfy. For instance, if the goal is to ensure the safety of LLM outputs, relevant properties might include that the LLM must not generate offensive language (Property 1) and must not encourage harmful behavior (Property 2). These properties form the basis for designing oracles that evaluate whether the SUT meets the defined goals.
2.3. Oracles
Oracles determine whether the SUT meets the defined properties we have decided upon given the test case goal. We make a novel distinction between the two levels of Atomic and Aggregated Oracles, and emphasize the latter’s importance in LLM-based testing.
An Atomic Oracle evaluates correctness based on a single test execution, much like in traditional software testing, where outputs are compared against predefined criteria. Atomic oracles may use strict deterministic checks, such as equality comparisons, regular expressions, or structural matching–effective for predictable outputs like numerical computations. However, LLM-generated responses introduce inherent variability, frequently making such rigid criteria insufficient. To address this, Atomic Oracles can also incorporate probabilistic or heuristic-based evaluations, including e.g. semantic similarity measures, rule-based checks for required keywords, and human-in-the-loop validation for subjective assessments. While these alternatives can be useful, they often will still fail to fully account for LLMs’ non-deterministic nature, even when the LLMs are configured for determinism (e.g., temperature = 0) (Atil et al., 2024).
An Aggregated Oracle mitigates this limitation by evaluating correctness across multiple test runs under the same conditions. Since LLM outputs vary for the same input, multiple samples can help establish consistency, and correctness is better assessed through statistical aggregation rather than single-instance evaluations. These oracles rely on Atomic Oracles for individual test executions and then apply an aggregation function to derive a final verdict. Aggregation strategies include measuring result variance, majority voting, confidence-weighted scoring, or reference-based comparisons against paraphrased expected outputs.
This distinction has major implications for LLM-based testing. Traditional methods struggle with non-deterministic outputs, necessitating batch testing frameworks that repeatedly execute test cases to compute aggregated verdicts. The choice of aggregation strategy depends on the application: structured data extraction may allow stricter Atomic Oracles, while creative text generation requires a more nuanced Aggregated Oracle approach.
By bridging deterministic software testing with LLM stochasticity, Aggregated Oracles provide a robust and realistic methodology for evaluating correctness in an inherently variable domain.
2.4. Inputs
Inputs are the concrete prompts that drive SUT execution, forming the final dimension of the taxonomy. They typically populate placeholders in prompt templates with specific dataset values, simulating real user interactions or system executions. However, in systems allowing multi-turn dialogues, having reasoning agents, or stateful architectures, later inputs often depend on earlier outputs, which can make static substitutions insufficient.
In addition to specifying the specific data values or how to derive sequences of inputs, the Inputs dimension can also list explicit variability operators to assess the SUT’s robustness. These fall into two categories: syntactic variations, which modify formatting or wording while preserving meaning, and semantic variations, which shift the meaning to explore different behaviors. Such variations apply at both the prompt template and input levels, expanding test coverage and blurring the distinction between prompt templates as input (from the LLM’s perspective) and as part of the SUT (as a development artefact).
This layered approach allows comprehensive probing of the SUT, generating diverse outputs for evaluation by oracles. Unlike traditional software testing, where inputs map directly to expected outputs, LLM interactions introduce ambiguity–user inputs may refine, complement, or complete a prompt, complicating clear test case boundaries.
Addressing input variability requires systematic coverage strategies across syntactic and semantic dimensions. This enhances reliability and robustness in LLM-based systems while clarifying failure points linked to input ambiguities.
Systematically addressing input variability involves creating comprehensive input coverage strategies that incorporate both syntactical and semantical dimensions. This approach not only ensures higher reliability and robustness of LLM-based software but also supports clearer identification of failure points related to input ambiguities.
3. Mapping the Taxonomy
Here we analyze how each facet of our taxonomy is addressed by existing tools that represent the state of practice in LLM application testing. Several open-source (Team, 2024a, b; Parea AI, 2025; Trulens Team, 2025; Comet, 2025; Giskard Team, 2024) and hosted (Weights & Biases, 2025; LangChain, 2025) platforms have emerged to support this domain. This paper focuses on a selection of widely used open-source frameworks as representative examples.
We identified relevant GitHub repositories by searching for “LLM Testing” and “LLM Evaluation”, selecting frameworks that specifically target LLM applications (not foundational models) and provide dedicated testing interfaces. The three selected projects, chosen based on GitHub popularity (stars) and distinct testing features, are Promptfoo (Team, 2024b) (5.6k stars), DeepEval (Team, 2024a) (5.1k stars), and Giskard (Giskard Team, 2024) (4.3k stars).
Table 1. Mapping of a selection of open-source testing frameworks to the taxonomy
Facets Promptfoo DeepEval Giskard
SUT SUT variations can be compared using visualization features. SUT variations across versions can be analyzed using visualization features to detect regressions. SUT variations are less emphasized but can be still compared via external platform integration.
Goal Testing goal can be set with one or more properties with varying number of inputs. Testing goal can be set with one or more properties for each test input. Testing goal is defined as one or more properties across a set of inputs.
Oracles Oracles are implemented using deterministic and model-based predefined metrics, allowing weighted metric combinations. Oracles are handled mostly using model-based metrics, including conversational/multi-modal support. Oracles are typically implemented as dataset-wise scores via built-in metric functions across multiple inputs.
Inputs Inputs with semantic variations can be synthesized with an separate LLM agent. Inputs with semantic variations can be synthesized with an separate LLM agent and with various resources (e.g., documents). Inputs can be transformed using various predefined syntactic transformation methods.
SUT
All tools accommodate variability across the three SUT sub-facets, though their capabilities differ slightly. Each allows users to specify the Component under test (e.g., a prompt template or agent), select the Model(s) (e.g., GPT-4o, Claude 3.7 Sonnet), and adjust Configurations (e.g., temperature, token limits) that influence runtime behavior. Notably, Promptfoo enables “unfolded” comparisons, allowing testers to assess differences between prompt versions, model choices, or configuration settings. Similarly, DeepEval provides visualization tools but primarily focuses on detecting regressions over time. For cross-version SUT comparisons, Giskard integrates with external commercial platforms like Weights and Biases (Weights & Biases, 2025).
Assessment: Each tool, to some degree, recognizes that developers frequently re-evaluate multiple implementations throughout development. They support continuous testing by enabling iterative modifications to components, models, and configurations, while also facilitating some level of structured comparisons of test results.
Goal
An ideal test case is defined with a clear and specific testing goal, typically composed of one or more relevant properties. In Promptfoo and DeepEval, a test case consists of a single input paired with an expected output, evaluated against one or more property-checking metrics. In contrast, Giskard integrates into an ML evaluation pipeline, running tests on entire datasets and producing dataset-wide scores. Some LLM testing sub-goals–such as assessing robustness against prompt injection attacks–are best evaluated using a set of targeted inputs. While Promptfoo and DeepEval can execute multiple test cases in a single evaluation, they lack explicit support for linking evaluations to a coherent test goal. Conversely, Giskard lacks the flexibility to configure individual test cases separately when needed.
Together, these frameworks fall short of balancing flexibility (checking properties at the individual input level) with coherence (grouping relevant inputs under a common evaluation goal).
Assessment: While these tools offer various property-checking mechanisms, they struggle with goal variability and representability–the ability to flexibly define and assess test goals at different levels of granularity. Although properties serve as concrete criteria derived from test goals, current frameworks do not explicitly model this distinction or integrate it into evaluation workflows. Crucially, no standardized approach differentiates goals from properties at the test case level.
Oracles
The tools support a range of oracles, often referred to as “metrics” or “assertions,” including deterministic checks (e.g., substring matching) and heuristic, model-based evaluations. Users can define custom metrics using tailored prompts for a separate LLM instance (LLM-as-a-judge(Gu et al., 2024; Zheng et al., 2023) or through code-based implementations. Web interfaces enable manual review and labeling, but human-in-the-loop validation remains underdeveloped. While tools allow human-provided labels, they lack mechanisms for iterative collaboration between humans and model evaluators.
Evaluations typically yield either a continuous score (e.g., similarity to a ground-truth output on a 0-1 scale) or a binary pass/fail result, with continuous scores often thresholded into binary outcomes. Giskard operates at the dataset level, aggregating individual assessments (e.g., reporting a 70% pass rate for politeness). However, LLM outputs can vary across repeated runs, and current tools lack aggregation methods to handle multiple evaluations of the same input within the same system under test (SUT). Instead, each run is treated as an independent outcome, ignoring stochastic variability.
We argue for aggregated oracles that account for repeated evaluations, using different aggregation strategies depending on the property under test and the source of variation. For example, a strict requirement such as “the response MUST follow the specified JSON format” should pass only if no violations occur across multiple runs. Conversely, a softer requirement like “the response should be polite” could be satisfied if the majority of runs meet the criterion.
Assessment: While these tools offer diverse oracle definitions, model-based heuristic oracles remain unstable, and deterministic oracles lack expressiveness. Existing dataset-level aggregation is limited, and current tools fail to systematically capture variability across repeated runs. A principled approach to aggregated oracles is needed, along with flexible aggregation schemes (e.g., majority vote, strict pass, confidence-based thresholds) tailored to different evaluation properties.
Inputs
Developers often struggle with a lack of data or inputs when testing LLM applications. Existing tools mitigate this by generating synthetic inputs using separate LLM instances or modifying existing inputs with predefined operators. These transformations span both syntactic (e.g., punctuation removal, case conversion, language change) and semantic (e.g., content diversification while preserving format) variations. Many tools also support adversarial input generation for security and safety testing, aligning with known attack methods (OWASP, 2025), such as injecting jail-breaking prompts or embedding encrypted forbidden content to bypass security filters.
Giskard extends syntactic transformations with metamorphic testing, ensuring that minor syntactic changes do not significantly alter outputs. However, these input generation methods lack well-defined quantitative objectives, making it difficult to ensure comprehensive test coverage. This raises key questions: How can we determine whether available inputs are sufficient for thorough testing? How can we optimize regression testing to balance cost and rigor?
Assessment: A major limitation in existing tools is the absence of clear test adequacy criteria for input variability. While they provide input synthesis and data collection features, they do not systematically assess whether variations sufficiently test SUT robustness.
4. Towards 2030
While our taxonomy provides a foundation for structured LLM testing, several critical areas require further research. A key direction is the development of a stepwise methodology for LLM test case design. Extending the taxonomy into a structured method would help practitioners systematically define, implement, and validate test cases, bridging the gap between conceptual classification and real-world application. In particular, this could clarify the transition from abstract test goals to atomic and aggregated oracles via SUT properties.
Testing is increasingly entangled with development in LLM-based systems, warranting further exploration of ideas from test-driven design. Traditionally a validation step, testing in LLMs now plays a core role throughout development. Insights from continuous testing paradigms could help integrate testing seamlessly into LLM evolution and maintenance.
Aggregation and automation of oracles remain key challenges in LLM testing. Guidance is neeed on selecting aggregation types, statistical thresholds, voting mechanisms, or adaptive criteria to balance correctness and efficiency. AI-driven test frameworks could enhance this by dynamically proposing atomic vs. aggregated oracles, tailoring selection to contextual requirements. Future work should explore how property-based testing (Claessen and Hughes, 2011) and statistical early stopping methods (Wald and Wolfowitz, 1948) can improve oracle robustness, interpretability, and efficiency while reducing human intervention. Developing adaptive selection strategies would further refine test automation, ensuring more reliable and scalable evaluation processes.
SUT variability complicates test stability. Just as environmental shifts introduce regressions in traditional software, LLM-based systems are sensitive to model as well data variations. While our taxonomy acknowledges this, structured approaches to modeling these effects remain underexplored. Empirical studies and tool support are needed to analyze, mitigate, and predict the impact of configuration changes, including their interaction with input variation strategies.
Finally, stronger collaboration between academia and industry is essential for advancing LLM testing. As AI-driven software evolves at an unprecedented pace, research must not only keep up but also lead in defining rigorous, scalable, and adaptive testing methodologies. Academia plays a crucial role in transforming industry insights into generalizable principles, grounding them in software testing theory, and evaluating their actual and long-term impact. Open-source frameworks like DeepEval, Promptfoo, and Giskard provide a foundation for standardization, but ensuring they evolve to meet future challenges requires active engagement from researchers, practitioners, and policymakers. By working together, we can build the testing strategies, methods, and tools needed for the next generation of LLM-based and hybrid software/AI systems.